OpenStudy (anonymous):

of 350 males athletes at a high school, some play only basketball, some play only baseball, and some do both. if 250 of males play basketball, and 120 play both sports how many of the males play baseball?

OpenStudy (anonymous):

Hold on a sec, I need to do something

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

I did this beginning of last year, so I'm trying to remember how to do it

OpenStudy (anonymous):

ok :P

OpenStudy (anonymous):

Okay, I remember. So subtract 120 from 250 to get how many people only play basketball which equals 130 then, 350 = 120 + 130 + x x being the amount that only plays baseball 350 = 250 + x 350 = 250 = x 100 = x there are 100 who only play baseball

OpenStudy (anonymous):

Ok thanks (:

OpenStudy (anonymous):

No problem, sorry it took me so long

OpenStudy (anonymous):

its ok